Predictive Theory for the Grain Boundary Character Distribution

K. Barmak katayun@andrew.cmu.edu

Department of Materials Science and Engineering

Carnegie Mellon University Pittsburgh, PA 15213

E. Eggeling eva.eggeling@fraunhofer.at

Fraunhofer Institut Computer Graphics Research IGD

Project Group Graz, A-8010 Austria

M. Emelianenko memelian@gmu.edu

Department of Mathematical Sciences

George Mason University Fairfax, VA 22030

Y. Epshteyn rina10@andrew.cmu.edu

D. Kinderlehrer davidk@cmu.edu

R. Sharp sharp@andrew.cmu.edu

S. Ta’asan shlomo@andrew.cmu.edu

Department of Mathematical Sciences

Carnegie Mellon University Pittsburgh, PA 15213